Please help me with this questiob,
Jo is testing a computer game. A character runs across the screen at a rate r given by the equation r=30t, where t is the number of seconds after the game begins. The rate r is in pixels per second.
a. Write an equation to represent the distance d the character travels, in pixels, after t seconds.
b. A computer monitor measures 12 inches across, and has a resolutions of 85 pixels per inch. About how long does it take the character to run across the screen? Round to the nearest tenth of a second.
Thank you in advance...

Answer by Theo(13342)

the figure travels at a rate of 30 pixels per second.

the formula would be distance = rate * time.

the distance would be equal to 12 inches * 85 pixels per inch = 1020 pixels.

the formula becomes 1020 pixels = 30 pixels per second * time.

solve for time to get time = (1020 pixels) / (30 pixels per second) = 34 seconds.

in 34 seconds, the character will run at a rate of 30 pixels per second and will achieve a distance of 34 * 30 = 1020 pixels which is a distance of 12 inches assuming 85 pixels per inch.

algebraically, you would probably solve as follows:

distance = rate * time.

d = distance
r = rate
t = time.

formula becomes d = r*t

d is given in pixels.
rate is given in pixels per second.
time is in seconds.

d = 12 inches * 85 pixels per inch = 1020 pixels.
r = 30 pixels per second.
t = time in seconds.

formula becomes 1020 = 30t

solve for t to get t = 1020/30 = 34 seconds.

unless there's something else going on that i don't know about, i believe this is what you are looking for.